OpenStudy (anonymous):

Find the absolute (global) Maximum and absolute (global) Minimum values of the function f(x)= (x^2)/(x^2-9) on the interval [-2,1].

OpenStudy (anonymous):

check \(f(-2), f(1)\) and then find the critical point by finding the derivative, set it equal to zero and solve check the function evaluated at that number as well

OpenStudy (anonymous):

f(-2)= -4/5 f(1)= -1/8.. the derivative is (2x(x^2-9)-x^2(2x))/(x^2-9)^2

OpenStudy (anonymous):

-18x/(x^2-9)^2=0

OpenStudy (anonymous):

x=0

OpenStudy (campbell_st):

an easy solution to this problem is to use some graphing software and graph it...